# Off-Policy Evaluation for Sequential Persuasion Process with Unobserved Confounding

### v1

@@ I Introduction
Strategic information sharing plays a critical role in economic interactions, policy design, and multi-agent systems[1, 2, 3]. Bayesian persuasion was first introduced by Kamenica and Gentzkow [4] as a powerful framework for analyzing how a sender can strategically reveal information to influence a receiver’s decisions. In the standard setting, a sender commits to an information disclosure policy before observing the state of the world, and the receiver, after observing the sender’s message, forms posterior beliefs and takes an action that affects both the sender’s and the receiver’s utilities.
Despite its theoretical elegance, Bayesian persuasion rests on assumptions that may not hold in practical settings. First, the framework presupposes that the sender possesses complete information about the receiver, including their observation process and all features that influence their decision-making (including utility functions). Second, it assumes that the receiver has sufficient knowledge about the underlying state space to make Bayesian inferences based on the sender’s signaling policy and the shared signal. Many research efforts have explored various directions to relax these restrictive assumptions [5, 6, 7] or to resolve computational challenges [8]. For example, Castiglioni et al. [9] introduced online Bayesian persuasion to address scenarios where the sender lacks knowledge of the receiver’s utility function, proposing an online learning approach to acquire this information iteratively. Several researchers have extended the framework to sequential decision-making processes. Gan et al. [10] and Wu et al. [11] formulated Bayesian persuasion within Markov decision processes, where a sender engages in multiple rounds of interaction with different myopic receivers at each timestep. Building on this foundation, Bacchiochchi et al. [12] leveraged reinforcement learning techniques for settings where the sender has limited prior knowledge of the environment. In a related vein, Lin et al. [13] developed the Markov signaling game framework and derived signaling gradients to facilitate reinforcement learning approaches.
@@
In this paper, we address this gap by extending the Bayesian persuasion framework to account for unobserved confounding in sender-receiver interactions. First, we introduce a formulation that models sequential persuasion with unobserved confounding as a Partially Observable Markov Decision Process (POMDP). This formulation captures the sender’s incomplete information about both the receiver’s belief dynamics and the unobserved confounder. We prove that finding an optimal observation-based policy in this POMDP is equivalent to solving for an optimal signaling strategy in the original persuasion framework. Second, we demonstrate how this reformulation enables the application of proximal learning, a causal inference technique introduced by Miao et al.[14] and extended further to POMDP settings by [15]. Using proximal learning, we address the issue of unobserved confounding for off-policy evaluation (OPE) in the persuasion process. This allows the sender to assess alternative signaling strategies using only observational data from a behavioral policy without requiring costly new experiments. Our approach provides a principled way to design robust persuasion strategies in the presence of confounding, with applications in recommendation systems, strategic communication, and human-AI interaction[16, 17].
The remainder of this paper is organized as follows. In Section II, we review the standard Bayesian persuasion framework, followed by a visualization of the causal graph of this framework. In Section III, we present our extension of the persuasion framework, formulating a sequential decision-making problem for the sender, and highlights the role of hidden confounders. In Section IV, we reformulate this problem as a POMDP and establish its equivalence to the persuasion framework. In Section V, we demonstrate how proximal learning can be applied to the POMDP to enable OPE for the persuasion framework. In Section VI, we present an application that can be modeled using the persuasion framework. Finally, in Section VII, we draw concluding remarks and highlight some future directions.
@@ II Bayesian Persuasion and Causality
Bayesian persuasion models an interaction between two agents, where a sender communicates using signals with a receiver who interacts with the environment based on this communication. Generally, the sender and receiver are noncooperative and have different objectives. Given that the receiver is self-interested and the sender knows more about the environment, the goal of the sender is to select a communication strategy to influence the receiver’s behavior in its favor. We illustrate the layout of this framework in Fig. 1.
@@
Let the state of the environment be denoted by the random variable $S$, and let any realization $s$ belong to a finite set $\mathcal{S}$. The state of the system is sampled from a distribution $\mu\in\Delta(\mathcal{S})$, also known as prior. In general, both agents know the prior $\mu$, while only the sender observes the realized state $s$. The receiver interacts with the environment by taking actions, denoted by the random variable $A$, with realization $a$ taking values from a finite set $\mathcal{A}$, and this interaction gives a reward to each of the agents. The sender and receiver obtain rewards $P^{s}$ and $P^{r}$ given by the function $\rho^{s}:\mathcal{S}\times\mathcal{A}\to[m^{s},M^{s}]$ and $\rho^{r}:\mathcal{S}\times\mathcal{A}\to[m^{r},M^{r}]$, respectively, where $m^{s}$, $M^{s}$, $m^{r}$, $M^{r}\in\mathbb{R}$ are the lower and upper bounds of their rewards.
In Bayesian persuasion, the sender can select a signaling policy $\pi$ from a finite set of policies denoted by $\mathcal{P}$. The sender then commits this signaling policy, $\pi:\mathcal{S}\to\Delta(\mathcal{Q})$, where $\mathcal{Q}$ is a finite space of signals. During the interaction, the sender samples a signal from signaling policy, i.e., $q\sim\pi(\cdot|s)$, and sends it to the receiver. Since the signal $q$ encodes information about the state $s$, the signal influences the receiver’s belief about the state $s$, which, in turn, will affect the self-interested action $a$.
Based on the signal, the receiver performs a Bayesian update of their belief about $S$ considering both the prior $\mu$ and the signaling policy $\pi$. For each $s\in\mathcal{S}$, the resulting posterior distribution is given by
@@
Then, the probability of taking action $A$ given the signal $q$ becomes deterministic, that is, $p^{\pi}(s|q)=1$ if $a=a^{*}(q;\pi)$ and $p^{\pi}(a|q)=0$, otherwise. We define the performance of a signaling policy $\pi$ for the sender as
$J(\pi)=\;$ $\mathbb{E}^{\pi}[\rho^{s}(S,A)],$ (3)
=\;= $\sum_{s\in\mathcal{S}}\sum_{q\in\mathcal{Q}}\;\mu(s)\;\pi(q|s)\;\rho^{s}(s,a^{*}(q;\pi)).$ (4)
The goal of the sender is to compute the best signaling policy $\pi^{*}$, i.e.,
$\pi^{*}=\arg\max_{\pi}J(\pi).$ (5)
Next, we discuss the causal graph of the Bayesian persuasion framework (see Fig. 2). This graph is made up of nodes and directed edges. Nodes denote random variables, which include the state, action, and rewards, which are the same quantities in the persuasion process. The posterior, which represents the belief of the receiver about the state of the system, is denoted by the variable $B$. The policy of the sender is indicated by $\Pi$. The edges, also known as causal links, indicate the direction of causality. Since the signal $Q$ is sampled from the sender’s signaling policy conditioned on the state, there exists a causal link from $S$ and $\Pi$ to the node $Q$. Furthermore, the belief update process is influenced by both the policy and the received signal, establishing causal links from these nodes to the belief $B$. The posterior, in turn, determines the receiver’s action $A$, leading to a directed edge from $B$ to $A$. Finally, by definition, both the state $S$ and the receiver’s action $A$, have direct causal links to the reward nodes $P^{r}$ and $P^{s}$.
@@ III Sequential Persuasion with unobserved confounding
In our framework of persuasion under unobserved confounding, we extend the standard Bayesian persuasion to a sequential setting where the sender communicates with the receiver and the receiver interacts with the environment over several rounds. The index of the round is indicated by $i$ and ranges from $i=0,\ldots,T$, where $T\in\mathbb{N}$ corresponds to the final round of interaction. The key distinction is the presence of an unobserved confounder that affects how the receiver updates their prior belief after communicating with the sender.
The state of the environment at each round $i$ is denoted by the random variable $S_{i}$, with realization $s_{i}\in\mathcal{S}$. Each state $S_{i}$ is sampled independently from the same distribution $\mu$. Unlike standard Bayesian persuasion, the true distribution $\mu$ is known only to the sender, not to the receiver. At each $i$, the sender can observe the state realization $s_{i}$. The sender can select a signaling policy $\pi_{i}$ from a finite set of policies denoted by $\mathcal{P}$ at each $i$. Then, the sender commits to a signaling policy $\pi_{i}$, samples signal $q_{i}\sim\pi_{i}(\cdot|s_{i})$, and shares them with the receiver. The receiver holds their own belief $B_{i}$ about state $S_{i}$, which is unknown to the sender. We consider that the receiver’s belief space is finite. Let this finite space of beliefs be denoted by $\mathscr{B}$. The receiver selects a self-interested action:
@@
where $b_{i}(s)$ is the belief $b_{i}$ based probability of a state realization $s$.
Remark 1.
In the context of human-AI recommendation systems, the AI communicates with a human who serves as the receiver. Typically, humans interpret and process information from these recommendations or signals in categorical terms rather than as nuanced probability distributions. For instance, an individual’s belief about the likelihood of rain on a given day may range from low to medium to high, reflecting their level of confidence. Additionally, due to computational constraints, a receiver may simplify their belief space by utilizing a finite set of categories, thus reducing the complexity involved in updating their beliefs to make self-interested decisions.
After the receiver selects an action, both the sender and the receiver receive rewards $P^{s}_{i}=\rho^{s}(S_{i},A_{i})$ and $P^{r}_{i}=\rho^{r}(S_{i},A_{i})$, respectively, which are bounded by $[m^{s},M^{s}]$ and $[m^{r},M^{r}]$. Let $\rho^{s}_{i}$ and $\rho^{r}_{i}$ denote reward realizations for $P^{s}_{i}$ and $P^{r}_{i}$, respectively. Since $\mathcal{S}$ and $\mathcal{A}$ are finite sets, we consider that $\rho^{s}_{i}$ and $\rho^{r}_{i}$ also belong to finite sets $\mathcal{R}^{s}$ and $\mathcal{R}^{r}$, respectively.
@@
The variable $Z$ represents the nature of the receiver. For instance, consider the conservativeness of a planner within a decision-making framework, which exists on a spectrum from risk-averse (pessimistic) to risk-neutral (optimistic). Similarly, the concept of trust in human-AI interactions varies from complete trust to total distrust. The inherent nature of the planner significantly influences their beliefs regarding the state of the system and their decision-making processes.
When the receiver obtains information about the signal and the committed policy, they update their belief based on an update rule given by the conditional distribution
$p(b_{i}|b_{i-1},\rho^{r}_{i-1},a_{i-1},q_{i},z_{i},\pi_{i}),$ (7)
where this belief-updating rule is unknown to the sender. Considering that the receiver has no information at the initial round, we assume that they start with a uniform prior; that is, $b_{0}$ is the uniform distribution on $\mathcal{Q}$. After the belief update, the receiver selects action $a^{*}_{i}$ according to (6). Since $Z_{i}$ is unobservable by the sender, and influences the receiver’s belief update, in the language of causal inference, $Z_{i}$ acts as an unobserved confounder. An unobserved confounder presents a challenge to the sender in keeping track of or predicting belief updates over the rounds. At each $i$, we denote all information available to the sender using the information vector, denoted by $\Delta_{i}$. Any realization of such information vector is given by a tuple $\delta_{i}=(s_{0:i-1},\pi_{0:i-1},q_{0:i-1},a_{0:i-1},\rho^{r}_{0:i-1})$. At each $i$, the set of possible realizations of $\Delta_{i}$ is denoted by $\mathcal{D}_{i}$.
@@
Next, we define the performance of any meta-policy $\bm{\pi}$ as
$J(\bm{\pi}):=\;$ $\mathbb{E}^{\bm{\pi}}[\sum_{i=0}^{T}\rho^{s}(S_{i},A_{i})],$ (8)
where $\mathbb{E}^{\bm{\pi}}(\cdot)$ is the expectation over the joint distribution induced by the meta-policy $\pi$ and the belief update in this expectation is given by (7) along with the self-interested receiver action selected according to (6).
@@
The goal of the sender is to compute the best signaling meta-policy $\bm{\pi}^{*}$, i.e.,
$\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\bm{\Pi}}J(\bm{\pi}).$ (9)
@@ IV Construction of POMDP Structure
This section demonstrates how the sequential persuasion framework with unobserved confounding can be formulated as a POMDP. Following the structure of a POMDP, the sender serves as an agent who makes a decision, the receiver becomes part of an environment, and the sender’s incomplete knowledge of the receiver’s belief and the unobserved confounder are considered unobserved quantities. We then consider a class of observation-based control strategies for the planner and formulate an optimization problem to find the best such strategy. We show that solving this problem is equivalent to solving Problem 1.
We reformulate the sequential persuasion process (SPP) as a system that evolves over discrete time steps, denoted by $t=0,\dots,T+1$, where $T$ is the final round of interaction in the persuasion process. Later, the definition of the state of this system will give intuition behind a horizon of $T+1$ time steps. To show that this system is a POMDP, we first construct the state and action at each $t$ based on the SPP and then prove that its evolution is Markovian. We use upper-case letters to denote random variables and lower-case letters to denote their corresponding realizations.
At each $t$, let $u_{t}$ denote the control action in this system. Starting with $i=t=0$, the state of the system $x_{0}$ is realized and the observation $y_{0}$ is available. Then, the agent implements a control action $u_{0}$, which is the sender’s signaling policy $\pi_{0}$ in the SPP. This concludes the first time step, after which the equivalent system evolves to the next time step. Although the equivalent system evolves, the round $i=0$ in the SPP is not complete. At the end of the time step $t=0$, only the state $s_{0}$ and the sender’s signal $\pi_{0}$ have been realized in the SPP. It is only after the equivalent system evolves to $t=1$ that the receiver in the SPP updates their belief in round $i=0$ and then selects action $a_{0}$ in the SPP. This is illustrated in Fig. 3.
@@
The state at time $t+1$ is given by $x_{t+1}=(\delta_{i+1},s_{i+1},b_{i},z_{i+1})$. Since $s_{i+1}\sim\mu$, it evolves independently and $z_{i+1}=z_{i}$ do not depend on the action at time $u_{t}=\pi_{i}$. The belief update, given by $p(b_{i}|b_{i-1},\rho^{r}_{i-1},a_{i-1},q_{i},z_{i},\pi_{i})$ depends on $(b_{i-1},\rho^{r}_{i-1},a_{i-1},q_{i},z_{i})$, which are all components of the state $x_{t}$ and the action $u_{t}=\pi_{i}$. Consider the information vector $\delta_{i+1}=(\delta_{i},s_{i},q_{i},a_{i},\rho^{r}_{i},\pi_{i})$. We know that $s_{i}\sim\mu$ is drawn independently. The receiver’s action $a_{i}$ is a function of the belief $b_{i}$, hence it is purely state and action-dependent. As the reward for the receiver $\rho^{s}_{i}$ is based on $a_{i}$ and $s_{i}$, it is also purely state and action-dependent. Given $s_{i}$, the action $\pi_{i}$ completely determines the signal as $q_{i}\sim\pi_{i}(\cdot|s_{i})$. This shows that the state $x_{t}$ and the action $u_{t}$ are enough to predict a distribution on the next state $x_{t+}$. Hence the state evolution is Markovian, that is, $p(x_{t+1}|x_{0:i},u_{0:i})=p(x_{t+1}|x_{t},u_{t})$. Since the agent cannot observe the full state $x_{t}$, this system is a POMDP. ∎
@@
We recall that in the persuasion framework, the sender’s meta-policy $\bm{\pi}$ defines a tuple of functional mappings $p_{i}:\mathcal{D}_{i}\times\mathcal{S}\to\mathcal{P}$. At each $i$, the signaling policy $\pi_{i}$ is set by evaluation of $p_{i}$ at $\Delta_{i}=\delta_{i}$ as $\pi_{i}=p_{i}(s_{0:i-1},\pi_{0:i-1},q_{0:i-1},a_{0:i-1}\rho^{r}_{0:i-1},q_{i})$. In the case of observation-based control strategies in the POMDP setting, the control law $g_{t}$ selects the action $u_{t}=\pi_{i}$ based on $u_{t}\;=g_{t}(y_{t})=g_{t}(s_{0:i-1},\pi_{0:i-1},q_{0:i-1},a_{0:i-1},\rho^{r}_{0:i-1},q_{i})$. With this, we observe that at each $t=i$, the selected signaling policy is identical in both frameworks. Consequently, the space of all sender meta-policies $\bm{\Pi}$ coincides with the set of all observation-based control strategies $\mathcal{G}$ for the planner in the POMDP framework.
@@
Since the set of all meta-policies $\bm{\Pi}$ coincides with the set of all observation-based control strategies $\mathcal{G}$, any strategy $\bm{g}\in\mathcal{G}$ is a valid meta-policy. From now on, we will use the notion of strategy in our analysis instead of meta-policy. Furthermore, we will use $u_{t}$, the agent’s action, to indicate the sender’s signaling policy.
@@
In the POMDP setting, we define the performance of any strategy $\bm{g}$ as
$J(\bm{g})=\mathbb{E}^{\bm{g}}\left[\sum_{t=0}^{T}r_{t}(X_{t+1})\right],$ (10)
where $\mathbb{E}^{\bm{g}}$ denotes the expectation on all the random variables with respect to the probability distributions generated by the choice of control strategy $\bm{g}$.
@@
The goal of the agent is to compute the best observation-based control strategy $\bm{g}^{*}$, i.e.,
$\bm{g}^{*}=\arg\max_{\bm{g}\in\bm{G}}J(\bm{g}).$ (11)
@@
Note that to formulate and solve Problem 1, we need to be able to compute the performance of any observation-based control strategy. At each $t$, let $\tau_{t}=(x_{0:t},y_{0:t},u_{0:t})$ be the trajectory of the system. We denote the space of all trajectories at time $t$ by $\mathcal{T}_{t}$. The observable trajectory of the system is denoted by $\tau^{o}_{t}=(y_{0:t},u_{0:t})$. The space of all observable trajectories at time $t$ is $\mathcal{T}^{o}_{t}$. Any strategy $\bm{g}$ induces a distribution $p^{\bm{g}}_{t}(\cdot)$ over $\mathcal{T}_{t}$. Similarly, it induces a distribution $p^{\bm{g},o}_{t}(\cdot)$ over $\mathcal{T}^{o}_{t}$. In Section V, we show that the performance of a strategy depends on the probability measure induced by this strategy on the trajectory space of the POMDP. In our analysis, we make use of the conditional independencies among variables in the POMDP based on their causal relationship. The causal relationships in the formulated POMDP are given by the causal graph illustrated in Fig. 4.

### v2

@@ I Introduction
Strategic information sharing plays a critical role in economic interactions, policy design, and multi-agent systems[1, 2, 3]. Bayesian persuasion (BP) was first introduced by Kamenica and Gentzkow [4] as a powerful framework for analyzing how a sender can strategically reveal information to influence a receiver’s decisions. In the standard setting, a sender commits to an information disclosure policy before observing the state of the world, and the receiver, after observing the sender’s message, forms posterior beliefs and takes an action that affects both the sender’s and the receiver’s utilities.
Despite its theoretical elegance, BP rests on assumptions that may not hold in practical settings. First, the framework presupposes that the sender possesses complete information about the receiver, including their observation process and all features that influence their decision-making (including utility functions). Second, it assumes that the receiver has sufficient knowledge about the underlying state space to make Bayesian inferences based on the sender’s signaling policy and the shared signal. Many research efforts have explored various directions to relax these restrictive assumptions [5, 6, 7] or to resolve computational challenges [8]. For example, Castiglioni et al. [9] introduced online BP to address scenarios where the sender lacks knowledge of the receiver’s utility function, proposing an online learning approach to acquire this information iteratively. Several works have extended the framework to sequential decision-making processes. Gan et al. [10] and Wu et al. [11] formulated BP within Markov decision processes, where a sender engages in multiple rounds of interaction with different myopic receivers at each timestep. Building on this foundation, Bacchiochchi et al. [12] leveraged reinforcement learning techniques for settings where the sender has limited prior knowledge of the environment. In a related vein, Lin et al. [13] developed the Markov signaling game framework and derived signaling gradients to facilitate reinforcement learning approaches.
@@
In this paper, we address this gap by extending the Bayesian persuasion framework to account for unobserved confounding in sender-receiver interactions. First, we introduce a formulation that models sequential persuasion with unobserved confounding as a Partially Observable Markov Decision Process (POMDP). This formulation captures the sender’s incomplete information about both the receiver’s belief dynamics and the unobserved confounder. We prove that finding an optimal observation-based policy in this POMDP is equivalent to solving for an optimal signaling strategy in the original persuasion framework. Second, we demonstrate how this reformulation enables the application of proximal learning, a causal inference technique introduced by Miao et al.[14] and extended further to POMDP settings by [15]. Using proximal learning, we address the issue of unobserved confounding for off-policy evaluation (OPE) in the persuasion process. This allows the sender to assess alternative signaling strategies using only observational data from a behavioral policy without requiring costly new experiments. The capability for OPE serves as a foundation for reinforcement learning through policy optimization. Our approach provides a principled way to design robust persuasion strategies in the presence of confounding, with applications in recommendation systems, strategic communications, and human-AI interaction[16, 17].
The remainder of this paper is organized as follows. In Section II, we review the standard Bayesian persuasion framework, followed by a visualization of the causal graph of this framework. In Section III, we present our extension of the persuasion framework, formulating a sequential decision-making problem for the sender, and highlights the role of hidden confounders. In Section IV, we formulate a POMDP, equivalent to the persuasion framework. In Section V, we demonstrate how proximal learning can be applied to the POMDP to enable OPE for the persuasion framework. In Section VI, we present an application that can be modeled using the persuasion framework. Finally, in Section VII, we draw concluding remarks and highlight some future directions.
@@ II Bayesian Persuasion and Causality
Bayesian persuasion illustrated in Fig. 1, models an interaction between two agents, where a sender communicates using signals with a receiver who interacts with the environment based on this communication.
@@
Let the state of the environment be denoted by the random variable $S$, and let any realization $s$ belong to a finite set $\mathcal{S}$. The state is sampled from a distribution $\mu\in\Delta(\mathcal{S})$, also known as prior. In BP, both agents know the prior $\mu$, while only the sender observes the realized state $s$. The receiver interacts with the environment by taking actions, denoted by the random variable $A$, with realization $a$ taking values from a finite set $\mathcal{A}$, and this interaction gives a reward to each of the agents. The sender and receiver obtain rewards $P^{s}$ and $P^{r}$ given by the function $\rho^{s}:\mathcal{S}\times\mathcal{A}\to[m^{s},M^{s}]$ and $\rho^{r}:\mathcal{S}\times\mathcal{A}\to[m^{r},M^{r}]$, respectively, where $m^{s}$, $M^{s}$, $m^{r}$, $M^{r}\in\mathbb{R}$ are the lower and upper bounds of their rewards.
In BP, the sender can select a signaling policy $\pi$ from a finite set of policies denoted by $\mathcal{P}$. The sender then commits this signaling policy, $\pi:\mathcal{S}\to\Delta(\mathcal{Q})$, where $\mathcal{Q}$ is a finite space of signals. During the interaction, the sender samples a signal from signaling policy, i.e., $q\sim\pi(\cdot|s)$, and sends it to the receiver. Based on the signal, the receiver performs a Bayesian update of their belief about $S$ considering both the prior $\mu$ and the signaling policy $\pi$. For each $s\in\mathcal{S}$, the resulting posterior distribution is given by
@@
Then, the probability of taking action $A$ given the signal $q$ becomes deterministic, that is, $p^{\pi}(a|q)=1$ if $a=a^{*}(q;\pi)$ and $p^{\pi}(a|q)=0$, otherwise. We define the performance of a signaling policy $\pi$ for the sender as
$J(\pi)=\sum_{s\in\mathcal{S}}\sum_{q\in\mathcal{Q}}\;\mu(s)\;\pi(q|s)\;\rho^{s}(s,a^{*}(q;\pi)).$ (3)
The goal of the sender is to compute the best signaling policy $\pi^{*}=\arg\max_{\pi}J(\pi)$.
Next, we discuss the causal graph of the BP framework (see Fig. 2). This graph is made up of nodes and directed edges. Nodes denote random variables, which include the state, action, and rewards, which are the same quantities in the persuasion process. The posterior, which represents the belief of the receiver about the state of the system, is denoted by the variable $B$. The policy of the sender is indicated by $\Pi$. The edges, also known as causal links, indicate the direction of causality. Since the signal $Q$ is sampled from the sender’s signaling policy conditioned on the state, there exists a causal link from $S$ and $\Pi$ to the node $Q$. Furthermore, the belief update process is influenced by both the policy and the received signal, establishing causal links from these nodes to the belief $B$. The posterior, in turn, determines the receiver’s action $A$, leading to a directed edge from $B$ to $A$. Finally, by definition, both the state $S$ and the receiver’s action $A$, have direct causal links to the reward nodes $P^{r}$ and $P^{s}$.
@@ III Sequential Persuasion with unobserved confounding
In our framework of persuasion under unobserved confounding, we extend the standard BP to a sequential setting where the sender communicates with the receiver and the receiver interacts with the environment over several rounds. The index of the round is indicated by $i$ and ranges from $i=0,\ldots,T$, where $T\in\mathbb{N}$ corresponds to the final round of interaction. The key distinction is the presence of an unobserved confounder that affects how the receiver updates their prior belief after communicating with the sender.
The state of the environment at each round $i$ is denoted by the random variable $S_{i}$, with realization $s_{i}\in\mathcal{S}$. Each state $S_{i}$ is sampled independently from the same distribution $\mu$. Unlike standard BP, the true distribution $\mu$ is known only to the sender, not to the receiver. At each $i$, the sender can observe $s_{i}$ and select a signaling policy $\pi_{i}$ from a finite set of policies denoted by $\mathcal{P}$ at each $i$. Then, the sender commits to a signaling policy $\pi_{i}$, samples signal $q_{i}\sim\pi_{i}(\cdot|s_{i})$, and shares them with the receiver. The receiver holds their own belief $B_{i}$ about state $S_{i}$, which is unknown to the sender. We consider that the receiver’s belief space is finite. Let this finite space of beliefs be denoted by $\mathscr{B}$. The receiver selects a self-interested action:
@@
where $b_{i}(s)$ is the belief $b_{i}$ based probability of a state realization $s$. After the receiver selects an action, both the sender and the receiver receive rewards $P^{s}_{i}=\rho^{s}(S_{i},A_{i})$ and $P^{r}_{i}=\rho^{r}(S_{i},A_{i})$, respectively, which are bounded by $m^{s},M^{s}]$ and $m^{r},M^{r}]$. Let $\rho^{s}_{i}$ and $\rho^{r}_{i}$ denote reward realizations for $P^{s}_{i}$ and $P^{r}_{i}$, respectively. Since $\mathcal{S}$ and $\mathcal{A}$ are finite sets, we consider that $\rho^{s}_{i}$ and $\rho^{r}_{i}$ also belong to finite sets $\mathcal{R}^{s}$ and $\mathcal{R}^{r}$, respectively.
@@
The variable $Z$ represents the nature of the receiver. For instance, consider the conservativeness of a planner within a decision-making framework, which exists on a spectrum from risk-averse (pessimistic) to risk-neutral (optimistic).
When the receiver obtains information about the signal and the committed policy, the dynamics fo their belief update is given by the conditional distribution $p(b_{i}|b_{i-1},\rho^{r}_{i-1},a_{i-1},q_{i},z_{i},\pi_{i})$, We note that the belief-update dynamics is unknown to the sender. Considering that the receiver has no information at the initial round, we assume that they start with a uniform prior; that is, $b_{0}$ is the uniform distribution on $\mathcal{S}$. After the belief update, the receiver selects action $a^{*}_{i}$ according to (4). Since $Z_{i}$ is unobservable by the sender, and influences the receiver’s belief update, in the language of causal inference, $Z_{i}$ acts as an unobserved confounder. An unobserved confounder presents a challenge to the sender in keeping track of or predicting belief updates over the rounds. At each $i$, we denote all information available to the sender using the information vector, denoted by $\Delta_{i}$. Any realization of such information vector is given by a tuple $\delta_{i}=(s_{0:i-1},\pi_{0:i-1},q_{0:i-1},a_{0:i-1},\rho^{r}_{0:i-1})$. At each $i$, the set of possible realizations of $\Delta_{i}$ is denoted by $\mathcal{D}_{i}$.
@@
Next, we define the performance of any meta-policy $\bm{\pi}$ as $J(\bm{\pi}):=\mathbb{E}^{\bm{\pi}}[\sum_{i=0}^{T}\rho^{s}(S_{i},A_{i})]$, where $\mathbb{E}^{\bm{\pi}}(\cdot)$ is the expectation over the joint distribution induced by the meta-policy $\bm{\pi}$, the belief update dynamics and the self-interested receiver action selected according to (4).
@@
The goal of the sender is to compute the best signaling meta-policy $\bm{\pi}^{*}$, defined as $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\bm{\Pi}}J(\bm{\pi})$.
@@ IV Construction of POMDP Structure
This section demonstrates how the sequential persuasion process (SPP) with unobserved confounding can be formulated as a POMDP. The sender serves as an agent who makes a decision, the receiver becomes part of an environment, and the sender’s incomplete knowledge of the receiver’s belief and the unobserved confounder are considered unobserved quantities. Then, we use this framework to formulate a problem equivalent to Problem 1.
We reformulate the SPP as a system that evolves over discrete time steps, denoted by $t=0,\dots,T+1$, where $T$ is the final round of interaction in the persuasion process. Furthermore, the definition of the state of this system will give intuition behind a horizon of $T+1$ time steps. In our analysis, we show how the newly formulated system is a valid POMDP. We use upper-case letters to denote random variables and lower-case letters to denote their corresponding realizations.
At each $t$, let $u_{t}$ denote the control action in this system. Starting with $i=t=0$, the state of the system $x_{0}$ is realized and the observation $y_{0}$ is available. The agent then implements a control action $u_{0}=\pi_{0}$, which is the signaling policy. This concludes the first time step, after which the equivalent system evolves to the next time step. Although the equivalent system evolves, the round $i=0$ in the SPP is not complete. At the end of the time step $t=0$, only the state $s_{0}$ and the sender’s signal $\pi_{0}$ have been realized in the SPP. It is only after the equivalent system evolves to $t=1$ that the receiver in the SPP updates their belief in round $i=0$ and then selects action $a_{0}$ in the SPP. This is illustrated in Fig. 3.
@@
The state at time $t+1$ is given by $x_{t+1}=(\delta_{i+1},s_{i+1},b_{i},z_{i+1})$. Since $s_{i+1}\sim\mu$, it evolves independently and $z_{i+1}=z_{i}$ do not depend on the action at time $u_{t}=\pi_{i}$. The belief update, given by $p(b_{i}|b_{i-1},\rho^{r}_{i-1},a_{i-1},q_{i},z_{i},\pi_{i})$ depends on $(b_{i-1},\rho^{r}_{i-1},a_{i-1},q_{i},z_{i})$, which are all components of the state $x_{t}$ and the action $u_{t}=\pi_{i}$. Consider the information vector $\delta_{i+1}=(\delta_{i},s_{i},q_{i},a_{i},\rho^{r}_{i},\pi_{i})$. We know that $s_{i}\sim\mu$ is drawn independently. The receiver’s action $a_{i}$ is a function of the belief $b_{i}$, hence it is purely state and action-dependent. As the reward for the receiver $\rho^{s}_{i}$ is based on $a_{i}$ and $s_{i}$, it is also purely state and action-dependent. Given $s_{i}$, the action $\pi_{i}$ completely determines the signal as $q_{i}\sim\pi_{i}(\cdot|s_{i})$. This shows that the state $x_{t}$ and the action $u_{t}$ are enough to predict a distribution on the next state $x_{t+1}$. Hence the state evolution is Markovian, that is, $p(x_{t+1}|x_{0:i},u_{0:i})=p(x_{t+1}|x_{t},u_{t})$. Since the agent cannot observe the full state $x_{t}$, this system is a POMDP. ∎
@@
We recall that in the persuasion framework, the sender’s meta-policy $\bm{\pi}$ defines a tuple of functional mappings $p_{i}:\mathcal{D}_{i}\times\mathcal{S}\to\mathcal{P}$. At each $i$, the signaling policy $\pi_{i}$ is set by evaluation of $p_{i}$ at $\Delta_{i}=\delta_{i}$ as $\pi_{i}=p_{i}(s_{0:i-1},\pi_{0:i-1},q_{0:i-1},a_{0:i-1},\rho^{r}_{0:i-1},q_{i})$. In the case of observation-based control strategies in the POMDP setting, the control law $g_{t}$ selects the action $u_{t}=\pi_{i}$ based on $u_{t}\;=g_{t}(y_{t})=g_{t}(s_{0:i-1},\pi_{0:i-1},q_{0:i-1},a_{0:i-1},\rho^{r}_{0:i-1},q_{i})$. With this, we observe that at each $t=i$, the selected signaling policy is identical in both frameworks. Consequently, the space of all sender meta-policies $\bm{\Pi}$ coincides with the set of all observation-based control strategies $\mathcal{G}$ for the planner in the POMDP framework.
@@
Since the set of all meta-policies $\bm{\Pi}$ coincides with the set of all observation-based control strategies $\mathcal{G}$, any strategy $\bm{g}\in\mathcal{G}$ is a valid meta-policy.
@@
In the POMDP setting, we define the performance of any strategy $\bm{g}$ as $J(\bm{g})=\mathbb{E}^{\bm{g}}\left[\sum_{t=0}^{T}r_{t}(X_{t+1})\right]$, where $\mathbb{E}^{\bm{g}}$ denotes the expectation on all the random variables with respect to the probability distributions generated by the choice of control strategy $\bm{g}$.
@@
The goal of the agent is to compute the best observation-based control strategy $\bm{g}^{*}$, defined as $\bm{g}^{*}=\arg\max_{\bm{g}\in\bm{G}}J(\bm{g})$.
@@
Note that to formulate and solve Problem 1, we need to be able to compute the performance of any observation-based control strategy. At each $t$, let $\tau_{t}=(x_{0:t},y_{0:t},u_{0:t})$ be the trajectory of the system. We denote the space of all trajectories at time $t$ by $\mathcal{T}_{t}$. The observable trajectory of the system is denoted by $\tau^{o}_{t}=(y_{0:t},u_{0:t})$. The space of all observable trajectories at time $t$ is $\mathcal{T}^{o}_{t}$. Any strategy $\bm{g}$ induces a distribution $p^{\bm{g}}_{t}(\cdot)$ over $\mathcal{T}_{t}$. Similarly, it induces a distribution $p^{\bm{g},o}_{t}(\cdot)$ over $\mathcal{T}^{o}_{t}$. In our analysis to compute the performance of any strategy, we make use of the conditional independencies among variables in the POMDP based on their causal relationship. The causal relationships in the formulated POMDP are given by the causal graph illustrated in Fig. 4.